OpenStudy (anonymous):

(3p over 5) + (8 over 5) = 1

OpenStudy (anonymous):

\[(3p/5)+(8p/5)=1 , 8p/5=1, 8p=5, p=5/8\]

OpenStudy (anonymous):

sure